3y-8x=-110
x-2y=30

Answer by mananth(16946)

-8 x + 3 y = -110 .............1
1 x -2 y = 30 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 3
-16 x 6 y = -220
3 x -6 y = 90
Add the two equations
-13 x = -130
/ -13
x = 10
plug value of x in (1)
-8 x + 3 y = -110
-80 + 3 y = -110
3 y = -110 + 80
3 y = -30
y = -10

m.ananth@hotmail.ca